# RLinf: Flexible and Efficient Large-scale Reinforcement Learning via Macro-to-Micro Flow Transformation

### v1

@@ 1 Introduction
The rapid progress of large language models (LLMs) has reached a point where further scaling model alone yields diminishing returns. To push intelligence beyond pretraining, reinforcement learning (RL) has emerged as a crucial paradigm. Recent advances such as RLHF [29, 5], GRPO [37], and RL for embodied agents [17, 18] and Deep Research [26, 50] all rely on RL to align LLMs with human preferences, improve reasoning, and enable autonomous interaction with complex environments. OpenAI and others predict that RL workloads will soon consume more computational resources than LLM pretraining [28], making RL training efficiency the most critical system concern.
However, efficient RL training for various scenarios such as reasoning, agentic and embodiment at the scale of modern large models is challenging, which combines highly heterogeneous components with diverse workload characteristics and resource demands, such as LLM generation, inference and training, reward models, critic models, agent tooling and embodied environment simulators. For instance, LLM training consumes more accelerator (e.g., GPU) memory than LLM generation and inference (prefill-only generation) to maintain gradients and optimizer states, while LLM generation shows high dynamicity in response lengths, leading to low accelerator utilization. Moreover, components like LLM training support diverse parallelization strategies (e.g., data, tensor, pipeline parallelism), whereas others scale only via instance replication and may yield computation workloads distinct from common tensor computation in LLM, e.g., embodied simulators [7, 41] that require CPU for physics simulation and GPU graphics pipeline for 3D rendering.
Single execution mode of existing RL training systems fails to capture this diversity, leading to suboptimal efficiency. Collocated execution, where components sequentially occupy accelerators [39], suffers from the long-tail problem due to varying generation lengths, leaving accelerators idle. Disaggregated pipelining, where components run concurrently on separate accelerators with pipelining [10], mitigates the long-tail issue but introduces memory and computation imbalance (§2.2). Neither mode is universally optimal. Many RL workloads demand hybrid scheduling of the components to maximize efficiency, i.e., mixing collocation and pipelining in a more flexible way. However, supporting such flexible execution modes for a single programmed workflow is a significant challenge, as they often require different program structures and communication patterns. Also, identifying the right scheduling for a given workflow usually requires considerable manual tunning.
@@ 2.1 RL Workflows in LLM Era
The simplest is GRPO [37], an RL algorithm designed to reduce reliance on reward models. It involves a single LLM that generates multiple responses, e.g., 8, for a query (i.e., Generation), computes logarithmic probabilities for these responses (i.e., Inference), and uses the results as training data to update the same model (i.e., Training). The updated weights are then synchronized back for inference and generation, completing one training iteration.
In contrast, the RLHF [29] workflow adopts PPO [36], resulting in a more complex design involving four LLMs. The actor model serves as the core policy, generating responses to queries. The reference model remains fixed to constrain the actor from drifting too far from its initialization. The reward model assigns scalar rewards to generated responses, while the critic model estimates expected rewards to stabilize training. Actor and critic are trainable, whereas reference and reward models are frozen. These components interact closely, as shown in the figure.
Beyond algorithms, RL workflows also vary by application scenario. In embodied intelligence [17, 18], RL relies on simulators that simulates the physical world. An LLM interacts with the simulator by generating actions and receiving feedback, producing trajectories that serve as training data. Similarly, in Deep Research [26, 50], RL improves model performance through interaction with a search server that retrieves online information. The resulting rollout results are fed into training, while inference follows the GRPO-style workflow.
@@ 2.2 Inefficiencies in Diverse RL Workflows
Dynamicity in Rollout Wastes Computation.   The rollout phase is inherently dynamic. Lengths fluctuate across responses of the same query, and even more so across different queries. Embodied tasks such as grasping can take different numbers of steps. In deep research, generating a report may involve varying numbers of search interactions. Since rollouts are executed in batches, these variations create a long-tail problem where a few slow queries block the entire phase from proceeding to inference or training. The problem commonly exists in the collocated mode (e.g., veRL [39]), where generation, inference, and training sequentially share GPUs by swapping between CPU and GPU memory. We conducted an experiment of a 7B math reasoning RL training [31] on 8 nodes with 8 H100 GPUs each, as shown in Figure 2. The number of unfinished responses quickly shrinks to less than 5%, where a very small set of long-tail responses stalls the generation, leaving many GPUs underutilized or idle. Scaling out with more GPUs worsens the problem as idle time grows.
@@ 3.1 Overview
To produce a desirable micro execution flow across these scheduling dimensions, the core of the scheduler module is a profiling-guided scheduling policy, which utilizes runtime profiling of worker characteristics to search for the optimal execution mode of each worker. Based on the determined execution mode, the Controller assigns workers to accelerators, manages inter-worker connections, and orchestrates the execution flow by dispatching function invocations to the corresponding workers. To realize this, two mechanisms named elastic pipelining and context switch are devised to enable spatial and temporal orchestration of workers, respectively. Adaptive communication utilities such as point-to-point communication and data channel act as the data plane to support scalable worker interactions. The entire system leverages Ray [24] to remotely launch and control workers.
@@ 3.2 Workflow Construction Interface
with out_channel.device_lock:
@@
The design philosophy of RLinf is to maximize system flexibility to achieve high efficiency, which is also the guideline for RLinf’s programming interface design. To this end, unlike traditional graph-based declarative programming [1] that sacrifices control flow flexibility, debuggability and transparency for optimization opportunity, RLinf adopts a procedural programming paradigm that enables developers to flexibly express workflows imperatively. An example workflow based on this interface is shown in Figure 5. As shown, an RL program based on RLinf consists of two parts: (1) worker programs that define the logic of each RL component (e.g., simulator, LLM generation, actor training), and (2) a workflow runner that orchestrates the overall workflow by invoking the workers’ core functions and defining inter-worker interactions.
@@ 3.3 M2Flow Transformation
Input : Workflow graph $G$, number of devices NN.
@@ 3 $T_{best},S_{best}\leftarrowFindSchedule$ ($G_{dag}$, NN, $D_{table}$);
11     return $T_{perf}$, $S_{node}$;
@@ 3.4 Scheduling Policy
RLinf offers a large scheduling space through flexible orchestration of workers. Finding the most efficient execution mode for a given workflow is challenging. Thus, RLinf introduces two modules: the profiler and the scheduler. The profiler measures each component’s execution time and memory usage under different granularity, i.e., the execution batch sizes. These results are then fed to the scheduler, which searches the space and determines the final placement and execution plan. Specifically, the plan specifies GPU assignments, pipelining configurations and data processing granularity.
To determine the optimal worker scheduling, RLinf constructs a workflow graph in a just-in-time manner. The graph is extracted during profiling, when the workflow is executed, by tracing the data flow among workers through the communication primitives.
@@
$T_{critical}+(M/m-1)\times T_{bottleneck},$
where $T_{critical}$ is the pipeline warmup and cooldown time, $T_{bottleneck}$ is the runtime of the slowest subgraph, $M$ is the total batch size, and $m$ is the data processing granularity.
@@ 3.5 Adaptive Communication
However, existing collective communication libraries for GPU and CPU data, e.g., NCCL [27], Gloo [30], and MPI [9], fail to meet our goals as they are mostly built for standard communication patterns across a fixed number of processes in traditional model training and serving scenarios. In contrast, RL components manifest considerable spatial and temporal dynamicity. Spatially, components can collocate on the same device, or being distributed across different devices. Temporally, components can be launched or terminated at arbitrary times. However, state-of-the-art libraries like NCCL does not support flexible rank scaling and efficient intra-device communications. Furthermore, the data communicated between components can be complex data structures beyond standard contiguous GPU/CPU data buffer, e.g., a composition of multiple data buffer with varying sizes. Efficient handling of such data dynamicity is also missing in existing libraries.
@@ 4 Implementation
Currently, RLinf supports not only traditional LLM-based reasoning RL, but also agentic RL with tooling, and embodied RL involving complex workloads like 3D rendering, physics simulation, and robotic control. For RL algorithms, RLinf supports popular algorithms like PPO [36], GRPO [37], DAPO [45], and REFINFORCE++ [11], as well as some of their off-policy asynchronous versions. For models, we have implemented support for language models like Qwen [31], multi-modal models like VLM [2], and embodied models like VLA [17, 16] and Pi0 [42]. The rich RL workflow, algorithm and model support not only accelerates the development of new RL workflows with RLinf, but also demonstrates its generality, versatility and extensibility in practice.
Cluster Management and Device Allocation.   RLinf leverages Ray to realize cluster management, launching worker processes on remote nodes, and dispatching worker function executions. RLinf does not rely on Ray for device allocation to workers, because Ray only supports rigid packed-style (i.e., consecutive devices) or spread-style (i.e., spread-across-node first) resource allocation [24], which are not flexible enough for today’s complex and dynamic RL workflows. Instead, RLinf offers a flexible device allocation strategy, which allows each worker process to be allocated with any device or devices of any node across the cluster, by simply specifying the target devices’ global IDs.
Failure Monitoring.   The multi-component and distributed nature of RLinf makes it prone to various types of failures, such as worker process crashes, node failures, and network issues. RLinf implements failure monitoring of worker processes by wrapping public worker function (which can be launched remotely) with a failure handler, which catches not only standard Python exceptions, but also registers signal handlers to catch native failures like segmentation faults and terminations. Upon catching any failure, the failure handler reports the failure, and commit suicide immediately to reduce log contamination of the failure chain effects. Meanwhile, the Controller runs a failure monitor thread for each worker process, which periodically checks the liveness of the worker process via Ray. Upon detecting a failure of any worker process, the Controller quickly kills the whole system, so as to avoid triggering misleading timeout errors on other workers requiring the output of the failed worker.
@@ 5 Evaluation
Figure 8: End-to-end throughput of RLinf and veRL under different cluster scale and model size settings.
Figure 9: End-to-end throughput of RLinf and baselines under different cluster scale and placement strategy.
5.1 Experimental Setup
Models.   We evaluate RLinf on two types of modern model families listed in Table 1. For reasoning LLMs, we choose Qwen2.5 Models [31] distilled by DeepSeek-R1 with sizes ranging from 1.5B to 32B, which are widely used in the research community. For embodied models, we choose OpenVLA [17] and OpenVLA-OFT [16] models, both supervised-finetuned by RL4VLA [21] and SimpleVLA-RL [18]. OpenVLA and OpenVLA-OFT are two models widely applied by embodiment research community.
Table 1: Experiments setting.
Model Environment Type Task Baseline
Qwen2.5-1.5B - LLM Text Generation veRL
Qwen2.5-7B - LLM Text Generation veRL
Qwen2.5-32B - LLM Text Generation veRL
OpenVLA ManiSkill VLA Embody RL4VLA
OpenVLA-OFT LIBERO VLA Embody SimpleVLA-RL
Datasets and Environments.   For reasoning LLMs, we use the AReaL-boba-Data dataset [43]. This dataset integrates multiple standard datasets including DeepScaleR, Open-Reasoner-Zero, Light-R1, DAPO, NuminaMath (AoPS/Olympiad subsets), and ZebraLogic. Overly simple problems are filtered out to ensure dataset quality and effectiveness.
For VLAs, we employ two training environments: the LIBERO [20] benchmark and a pick-and-place task formulated in RL4VLA [21], which is implemented on the ManiSkill [41] framework. We refer to the latter as ManiSkill for brevity in the following sections.
Configuration.   The model and environment configurations for LLM and embodied model experiments are detailed in Table 2 and Table 3, respectively. These settings remain consistent when scaling to a larger number of GPUs.
Table 2: Experiments setting of reasoning RL
Model Size 1.5B 7B 32B
Rollout Batch Size 512 512 512
Group Size 16 32 32
Sequence Length 28672 28672 28672
Actor TP Size 2 4 8
Rollout TP Size 1 2 4
Table 3: Experiments setting of embodied RL
Environment ManiSkill LIBERO
Number of Environments 256 512
Step 80 64
Testbed.   We deploy RLinf on a H100 cluster with 32 nodes and 256 GPUs. Each node has 8 NVIDIA H100-80GB GPUs, and 2 Intel Xeon Platinum 8558@2.1 GHz CPUs with 48 core and 2TB memory. Intra-node communication utilizes NVLink, while inter-node communication uses 8 Mellanox ConectX-7 RDMA NiCs per node, each provides 400 Gbps bandwidth with RoCEv2.
Baselines.   For LLMs, we compare RLinf against veRL [39] v0.5, the state-of-the-art open-source RLHF system. We choose SGLang [49] as its rollout engine and Megatron-LM [40] as its training backend.
For embodied AI models, we compare RLinf against SimpleVLA-RL [18] and RL4VLA [21]. SimpleVLA-RL is built upon the veRL framework. It extend veRL with parallel multi-environment rendering for faster sampling, and adapt it into an integrated training–inference–rendering framework. RL4VLA is a recipe for efficient PPO training on VLAs, we integrated its core algorithm into RLinf and applied key optimizations. This ensures a fair and efficient comparison.
Metrics.   For the LLM end-to-end experiment, we use RLHF throughput (tokens/sec) as the performance metric following [39], which is defined as dividing the total number of tokens in prompts and responses in a global batch by one RLHF iteration time. For the embodied models, we use batch throughput (batches/sec), computed by dividing the number of batches by iteration time in one step. All reported performance numbers are averaged over 10 training iterations after warm-up.
RL Algorithms.   For reasoning RL, we use the popular GRPO algorithm with following modifications:
@@ 5.1 Experimental Setup
    Token-Level Loss: Instead of averaging loss over the entire response sequence, we compute the average over tokens, as in DAPO. This prevents excessively long responses from dominating training and reduces their gradient impact.

### v2

@@ 1 Introduction
The rapid progress of large language models (LLMs) has reached a point where further scaling model alone yields diminishing returns. To push intelligence beyond pretraining, reinforcement learning (RL) has emerged as a crucial paradigm. Recent advances such as RLHF [rlhf, deeprl], GRPO [grpo], and RL for embodied agents [kim24openvla, simplevla] and Deep Research [nakano2021webgpt, zheng2025deepresearcher] all rely on RL to align LLMs with human preferences, improve reasoning, and enable autonomous interaction with complex environments. OpenAI and others predict that RL workloads will soon consume more computational resources than LLM pretraining [learningtoreasoning], making RL training efficiency the most critical system concern.
However, efficient RL training for various scenarios such as reasoning, agentic and embodiment at the scale of modern large models is challenging, which combines highly heterogeneous components with diverse workload characteristics and resource demands, such as LLM generation, inference and training, reward models, critic models, agent tooling and embodied environment simulators. For instance, LLM training consumes more accelerator (e.g., GPU) memory than LLM generation and inference (prefill-only generation) to maintain gradients and optimizer states, while LLM generation shows high dynamicity in response lengths, leading to low accelerator utilization. Moreover, components like LLM training support diverse parallelization strategies (e.g., data, tensor, pipeline parallelism), whereas others scale only via instance replication and may yield computation workloads distinct from common tensor computation in LLM, e.g., embodied simulators [duan2022survey, taomaniskill3] that require CPU for physics simulation and GPU graphics pipeline for 3D rendering.
Single execution mode of existing RL training systems fails to capture this diversity, leading to suboptimal efficiency. Collocated execution, where components sequentially occupy accelerators [verl], suffers from the long-tail problem due to varying generation lengths, leaving accelerators idle. Disaggregated pipelining, where components run concurrently on separate accelerators with pipelining [areal], mitigates the long-tail issue but introduces memory and computation imbalance (§2.2). Neither mode is universally optimal. Many RL workloads demand hybrid scheduling of the components to maximize efficiency, i.e., mixing collocation and pipelining in a more flexible way. However, supporting such flexible execution modes for a single programmed workflow is a significant challenge, as they often require different program structures and communication patterns. Also, identifying the right scheduling for a given workflow usually requires considerable manual tunning.
@@ 2.1 RL Workflows in LLM Era
The simplest is GRPO [grpo], an RL algorithm designed to reduce reliance on reward models. It involves a single LLM that generates multiple responses, e.g., 8, for a query (i.e., Generation), computes logarithmic probabilities for these responses (i.e., Inference), and uses the results as training data to update the same model (i.e., Training). The updated weights are then synchronized back for inference and generation, completing one training iteration.
In contrast, the RLHF [rlhf] workflow adopts PPO [ppo], resulting in a more complex design involving four LLMs. The actor model serves as the core policy, generating responses to queries. The reference model remains fixed to constrain the actor from drifting too far from its initialization. The reward model assigns scalar rewards to generated responses, while the critic model estimates expected rewards to stabilize training. Actor and critic are trainable, whereas reference and reward models are frozen. These components interact closely, as shown in the figure.
Beyond algorithms, RL workflows also vary by application scenario. In embodied intelligence [kim24openvla, simplevla], RL relies on simulators that simulates the physical world. An LLM interacts with the simulator by generating actions and receiving feedback, producing trajectories that serve as training data. Similarly, in Deep Research [nakano2021webgpt, zheng2025deepresearcher], RL improves model performance through interaction with a search server that retrieves online information. The resulting rollout results are fed into training, while inference follows the GRPO workflow.
@@ 2.2 Inefficiencies in Diverse RL Workflows
Dynamicity in Rollout Wastes Computation.   The rollout phase is inherently dynamic. Lengths fluctuate across responses of the same query, and even more so across different queries. Embodied tasks such as grasping can take different numbers of steps. In deep research, generating a report may involve varying numbers of search interactions. Since rollouts are executed in batches, these variations create a long-tail problem where a few slow queries block the entire phase from proceeding to inference or training. The problem commonly exists in the collocated mode (e.g., veRL [verl]), where generation, inference, and training sequentially share GPUs by swapping between CPU and GPU memory. We conducted an experiment of a 7B math reasoning RL training [qwen2025qwen25technicalreport] on 8 nodes with 8 H100 GPUs each, as shown in Figure 2. The number of unfinished responses quickly shrinks to less than 5%, where a very small set of long-tail responses stalls the generation, leaving many GPUs underutilized or idle. Scaling out with more GPUs worsens the problem as idle time grows.
@@ 3.1 Overview
To produce a desirable micro execution flow across these scheduling dimensions, the core of the scheduler module is a profiling-guided scheduling policy, which utilizes runtime profiling of worker characteristics to search for the optimal execution mode of each worker. Based on the determined execution mode, the Controller assigns workers to accelerators, manages inter-worker connections, and orchestrates the execution flow by dispatching function invocations to the corresponding workers. To realize this, two mechanisms named elastic pipelining and context switch are devised to enable spatial and temporal orchestration of workers, respectively. Adaptive communication utilities such as point-to-point communication and data channel act as the data plane to support scalable worker interactions. The entire system leverages Ray [ray] to remotely launch and control workers.
@@ 3.2 Workflow Construction Interface
with self.device_lock:
@@
The design philosophy of RLinf is to maximize system flexibility to achieve high efficiency, which is also the guideline for RLinf’s programming interface design. To this end, unlike traditional graph-based declarative programming [abadi2016tensorflow] that sacrifices control flow flexibility, debuggability and transparency for optimization opportunity, RLinf adopts a procedural programming paradigm that enables developers to flexibly express workflows imperatively. An example workflow based on this interface is shown in Figure 5. As shown, an RL program based on RLinf consists of two parts: (1) worker programs that define the logic of each RL component (e.g., simulator, LLM generation, actor training), and (2) a workflow runner that orchestrates the overall workflow by invoking the workers’ core functions and defining inter-worker interactions.
@@ 3.3 M2Flow Transformation
Input : Workflow graph $G$, execution time estimation function $E$ for each component, and number of devices NN.
@@ 3 $T_{best},S_{best}\leftarrowFindSchedule$ ($G_{dag}$, NN, $D_{table}$);
11     return $E_{node}$, $S_{node}$;
@@ 3.4 Scheduling Policy
RLinf offers a large scheduling space through flexible orchestration of workers, thus finding the most efficient execution mode is challenging. To this end, RLinf introduces two modules: the profiler and the scheduler. The profiler is used to measure and estimate the execution characteristics of each component under different numbers of GPUs. The scheduler then utilizes this information to composite an overall execution plan, containing specific GPU assignments and pipelining configurations.
Profiler.   The profiler measures each component’s execution time and memory usage under different data parallel sizes, because data parallel size tend to have positive (sometimes nearly linear) relation with the component throughput as data is consumed faster with the increasing data parallel size. For non-model components like simulators, the data parallel size is the number of instances. For model training and generation components, the profiler requires users to provide model parallel configuration to decide the data parallel size, which can usually be determined by the model size and GPU memory. With the profiled data, the profiler extrapolates the execution time and memory usage for larger data parallel sizes using polynomial extrapolation, outputting an execution time estimation function $E$ for each component. This function is then fed to the scheduler for it to estimate each component’s execution time and whether it can be fitted into memory given a selected number of GPUs. Moreover, during profiling, the overall RL workflow is also captured into a workflow graph in a just-in-time manner, by tracing the data flow among workers through the communication primitives.
@@
$T_{\text{critical}}+(M/m-1)\times T_{\text{bottleneck}},$
where $T_{\text{critical}}$ is the pipeline warm up and cool down time, $T_{\text{bottleneck}}$ is the runtime of the slowest subgraph, $M$ is the total batch size, and $m$ is the data processing granularity.
@@ 3.5 Adaptive Communication
However, existing collective communication libraries for GPU and CPU data, e.g., NCCL [nccl], Gloo [gloo], and MPI [mpi], fail to meet our goals as they are mostly built for standard communication patterns across a fixed number of processes in traditional model training and serving scenarios. In contrast, RL components manifest considerable spatial and temporal dynamicity. Spatially, components can collocate on the same device, or being distributed across different devices. Temporally, components can be launched or terminated at arbitrary times. However, state-of-the-art libraries like NCCL does not support flexible rank scaling and efficient intra-device communications. Furthermore, the data communicated between components can be complex data structures beyond standard contiguous GPU/CPU data buffer, e.g., a composition of multiple data buffer with varying sizes. Efficient handling of such data dynamicity is also missing in existing libraries.
@@ 4 Implementation
Currently, RLinf supports not only traditional LLM-based reasoning RL, but also agentic RL with tooling, and embodied RL involving complex workloads like 3D rendering, physics simulation, and robotic control. For RL algorithms, RLinf supports popular algorithms like PPO [ppo], GRPO [grpo], DAPO [yu2025dapo], and REINFORCE++ [hu2025reinforce++], as well as some of their off-policy asynchronous versions. For models, we have implemented support for language models like Qwen [qwen2025qwen25technicalreport], multi-modal models like VLM [bai2025qwen2], and embodied models like VLA [kim24openvla, kim2025fine] and Pi0 [openpi]. The rich RL workflow, algorithm and model support not only accelerates the development of new RL workflows with RLinf, but also demonstrates its generality, versatility and extensibility in practice.
Cluster Management and Device Allocation.   RLinf leverages Ray to realize cluster management, launching worker processes on remote nodes, and dispatching worker function executions. RLinf does not rely on Ray for device allocation to workers, because Ray only supports rigid packed-style (i.e., consecutive devices) or spread-style (i.e., spread-across-node first) resource allocation [ray], which are not flexible enough for today’s complex and dynamic RL workflows. Instead, RLinf offers a flexible device allocation strategy, which allows each worker process to be allocated with any device or devices of any node across the cluster, by simply specifying the target devices’ global IDs. Also, hardware devices beyond accelerators such as robot arms are also abstracted and managed as schedulable devices in the same manner as accelerators. This enables RLinf to scale on all kinds of hardware that foundation models can interact with.
@@ 5 Evaluation
We extensively evaluate RLinf across math-reasoning and embodied RL workloads, covering four different models of different sizes (i.e., Qwen2.5 [qwen2025qwen25technicalreport], Qwen3-MoE [yang2025qwen3], OpenVLA [kim24openvla], OpenVLA-OFT [kim2025fine]), two RL algorithms (i.e., GRPO [grpo], PPO [ppo]), and multiple cluster scales. Overall, our key findings include:
  • $\bullet$
    RLinf consistently outperforms state-of-the-art RL systems veRL [verl] and Slime [slime] by $\times\sim1.70$$\times$ on a variety of math-reasoning RL settings. The results also show that different RL settings favor different execution modes.
  • $\bullet$
@@
    The scheduling policy identifies the best execution mode within $7\times 10^{-4}\sim 5.98s$ on the clusters of 8 to 1024 GPUs.